### v2

# A note on Lie and Jordan structures of Leavitt path algebras

Abstract.
Let $L_{K}(E)$ be the Leavitt path algebra of a directed graph $E$ over a field $K$. In this paper, we determine $E$ and $K$ for the Lie algebra $\mathbf{K}_{L_{K}(E)}$ and the Jordan algebra $\mathbf{S}_{L_{K}(E)}$ arising from $L_{K}(E)$ with respect to the standard involution to be solvable.
Key words and phrases:
leavitt path algebra; Cohn path algebra; Lie algebra; Jordan algebra
2020 Mathematics Subject Classification. 16S88; 16W10
Corresponding author
1. Introduction
Every associative algebra $\mathcal{A}$ over a field $K$ can give rise to two new structures: the Lie algebra, denoted by $\mathcal{A}^{-}$, and the Jordan algebra, denoted by $\mathcal{A}^{\circ}$ over $K$. These structures can be obtained by defining new products within $\mathcal{A}$, namely the Lie bracket [x,y]=xyyx[x,y]=xy-yx for Lie algebras and the Jordan product $x\circ y=xy+yx$ for Jordan algebras respectively, for every $x,y\in\mathcal{A}$. A reasonable question arises: Are the properties of $\mathcal{A}$ as a ring mirrored in its Lie or Jordan algebra counterparts? Specifically, does the simplicity of $\mathcal{A}$ as a ring guarantee its simplicity as a Lie or Jordan algebra? In a seminal paper [11] by Herstein, this question was answered affirmatively. Assume further that $\mathcal{A}$ is equipped with an involution $\star$; that is, a map $^{\star}:\mathcal{A}\to\mathcal{A}$ satisfying the following conditions:
  1. (1)
    $(a+b)^{\star}=a^{\star}+b^{\star}$,
  2. (2)
    $(ab)^{\star}=b^{\star}a^{\star}$,
  3. (3)
    $(a^{\star})^{\star}=a$,
for all $a,b\in\mathcal{A}$. Let
$\mathbf{S}_{\mathcal{A}}=\{x\in\mathcal{A}:x^{\star}=x\}\text{ and }\mathbf{K}_{\mathcal{A}}=\{x\in\mathcal{A}:x^{\star}=-x\}$
be the sets of symmetric elements and skew-symmetric elements of $\mathcal{A}$ with respect to $\star$}, respectively. It is obvious that $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{A}}$ are Lie subalgebra and Jordan subalgebra of $\mathcal{A}^{-}$ and $\mathcal{A}^{\circ}$ respectively. These are the most important structures of Lie or Jordan algebras arising from an associative algebra. In particular, when $\mathcal{A}$ is the full matrix ring $\mathbb{M}_{n}(K)$ and is the transpose operation, then $\mathbf{K}_{\mathcal{A}}$ is identified with the Lie algebra of skew-symmetric matrices which corresponds to the Lie group of orthogonal matrices. Under certain assumptions on $\mathcal{A}$, Lie and Jordan structures of $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{A}}$ respectively were intensively investigated by Herstein in a series of papers ([7]-[11]) and by other authors in the references therein.
In this paper, we consider the case when $\mathcal{A}$ is the Leavitt path algebra $L_{K}(E)$ of a directed graph $E$ over a field $K$. The concept of the Leavitt path algebra $L_{K}(E)$ associated with a graph $E$ and coefficients from a field $K$ made its debut in [2], sparking widespread interest among scholars. These algebras were concurrently and separately introduced in [6], despite being published two years following [2]. Within the past few years, the Lie subalgebras of $L_{K}(E)^{-}$ arising from $L_{K}(E)$ have been investigated by many authors. Motivated by the groundbreaking work of Herstein [11], necessary and sufficient conditions for the Lie algebra $L_{K}(E),L_{K}(E)]$ of $L_{K}(E)^{-}$ to be simple were given in [3] and [4], under the assumption that $E$ is row-finite. In another direction, the paper [12] completely identified the graph $E$ and the field $K$ for the Lie algebra $L_{K}(E)^{-}$ to be solvable.
Let $E$ be a graph and $K$ a field. Then, the mapping $^{\star}:L_{K}(E)\to L_{K}(E)$ which sends $v$ to $v$, $e$ to $e^{*}$ and $e^{*}$ to $e$ for all $v\in E^{0}$ and $e\in E^{1}$, and leaves all elements in $K$ fixed, extends to an involution of $L_{K}(E)$. We call this the standard involution on $L_{K}(E)$ (see also [1, page 36]). Let $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ be the Lie subalgebra of $L_{K}(E)^{-}$ and the Jordan subalgebra of $L_{K}(E)^{\circ}$ with respect to this involution, respectively. In [5] the simplicity of the Lie algebra $\mathbf{K}_{L_{K}(E)}$ was also studied. The current paper is devoted to a study of the Lie algebra $\mathbf{K}_{L_{K}(E)}$ and the Jordan algebra $\mathbf{S}_{L_{K}(E)}$. To be more specific, we identify the graph $E$ and the field $K$ for the Lie algebra $\mathbf{K}_{L_{K}(E)}$ and the Jordan algebra $\mathbf{S}_{L_{K}(E)}$ to be solvable.
2. Preliminaries
2.1. Directed graphs and Leavitt path algebras.
Throughout this paper, the basic notations and conventions are taken from [1]. In particular, a (directed) graph $E=(E^{0},E^{1},r,s)$ consists of two sets $E^{0}$ and $E^{1}$ together with maps $r,s:E^{1}\to E^{0}$. The elements of $E^{0}$ are called vertices and the elements of $E^{1}$ edges. For the convenience of readers, in the following, we outline some concepts and notations that we use substantially in this paper.
Let $E$ be a graph. A vertex $v$ in $E$ is a sink if it emits no edges, while it is an infinite emitter if it emits infinitely many edges. A graph $E$ is said to be row-finite if $E$ contains no infinite emitter. A vertex $v$ is said to be regular if it is neither a sink nor an infinite emitter. A path $\mu$ in a graph $E$ is a finite sequence of edges $\mu=e_{1}e_{2}\dots e_{n}$ with $r(e_{i})=s(e_{i+1})$ for all $1\leq i\leq n-1$. We set $s(\mu):=s(e_{1})$ and $r(\mu):=r(e_{n})$. The set of all finite paths in $E$ is denoted by ${\rm Path}(E)$. An exit for a path $\mu=e_{1}\dots e_{n}$ is an edge $e$ such that $s(e)=s(e_{i})$ for some $i$ and $e\neq e_{i}$.
Definition 2.1 (Leavitt path algebra).
Let $E$ be a graph and $K$ a field. We define a set $\left(E^{1}\right)^{*}$ consisting of symbols of the form $\{e^{*}|e\in E^{1}\}$. The Leavitt path algebra of $E$ with coefficients in $K$, denoted by $L_{K}(E)$, is the free associative $K$-algebra generated by $E^{0}\cup E^{1}\cup(E^{1})^{*}$, subject to the following relations:
  1. (V)
    $vv^{\prime}=\delta_{v,v^{\prime}}v$ for all $v,v^{\prime}\in E^{0}$,
  2. (E1)
    $s(e)e=er(e)=e$ for all $e\in E^{1}$,
  3. (E2)
    $r(e)e^{*}=e^{*}s(e)=e^{*}$ for all $e\in E^{1}$,
  4. (CK1)
    $e^{*}f=\delta_{e,f}r(e)$ for all $e,f\in E^{1}$, and
  5. (CK2)
    $v=\sum_{\{e\in E^{1}|s(e)=v\}}ee^{*}$ for every $v\in{\rm Reg}(E)$.
Let $E$ be a graph. For each $e\in E^{1}$, we call $e^{*}$ the ghost edge. Also, the notation $\mu^{*}=e_{n}^{*}\dots e_{2}^{*}e_{1}^{*}$ stands for the corresponding ghost path of the path $\mu=e_{1}\dots e_{n}$.
2.2. Lie and Jordan structures of associative algebras
Every associative algebra $\mathcal{A}$ over a field $K$ is endowed with a Lie algebra structure for the Lie bracket [a,b]=abba[a,b]=ab-ba for all $a,b\in\mathcal{A}$. We denote this Lie algebra by $\mathcal{A}^{-}$. For $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we define $S,T]={\rm span}_{K}\{ab-ba:a\in S,b\in T\}$. A $K$-subspace $\mathcal{L}$ of $\mathcal{A}$ is called a Lie subalgebra of $\mathcal{A}^{-}$ if $a,b]\in\mathcal{L}$ for all $a,b\in\mathcal{L}$. For the Lie subalgebra $\mathcal{L}$ of $\mathcal{A}^{-}$, we define inductively the two series
$\mathcal{L}^{(0)}=\mathcal{L},\;\mathcal{L}^{(1)}=[\mathcal{L},\mathcal{L}],\;\dots,\;\mathcal{L}^{(n)}=[\mathcal{L}^{(n-1)},\mathcal{L}^{(n-1)}],\dots$
and
$\mathcal{L}^{0}=\mathcal{L},\;\mathcal{L}^{0}=[\mathcal{L},\mathcal{L}],\;\dots,\;\mathcal{L}^{n}=[\mathcal{L}^{n-1},L],\dots$
The Lie subalgebra $\mathcal{L}$ is called Lie solvable (resp., Lie nilpotent) if there exists $n\geq 0$ (resp., $m\geq 0$) such that $\mathcal{L}^{(n)}=0$ (resp., $\mathcal{L}^{m}=0$). If such a smallest $n$ (resp., $m$) exists, we will say that $\mathcal{L}$ is Lie solvable of index $n$ (resp., Lie nilpotent of index $m$).
Similarly, a Jordan product can be defined on $\mathcal{A}$ by $a\circ b=ab+ba$, for all $a,b\in\mathcal{A}$. If $\mathcal{A}$ is viewed as a Jordan algebra for the product $\circ$, we denote it by $\mathcal{A}^{\circ}$. Note that in the case ${\rm char}(K)=2$, the two structures $\mathcal{A}^{\circ}$ and $\mathcal{A}^{-}$ coincide. A $K$-subspace $\mathcal{J}$ of $\mathcal{A}$ is called a Jordan subalgebra of $\mathcal{A}^{\circ}$ if $a\circ b\in\mathcal{J}$ for all $a,b\in\mathcal{J}$. For two $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we define
$S\circ T={\rm span}_{K}\{ab+ba:a\in S,b\in T\}.$
For the Jordan subalgebra $\mathcal{J}$ of $\mathcal{A}^{\circ}$, we define the analogous series for $\mathcal{J}$ as follows:
$\mathcal{J}_{(0)}=\mathcal{J},\;\mathcal{J}_{(1)}=\mathcal{J}\circ\mathcal{J},\;\dots,\;\mathcal{J}_{(n)}=\mathcal{J}_{(n-1)}\circ\mathcal{J}_{(n-1)},\dots$
and
$\mathcal{J}_{0}=\mathcal{J},\;\mathcal{J}_{1}=\mathcal{J}\circ\mathcal{J},\;\dots,\;\mathcal{J}_{n}=\mathcal{J}_{n-1}\circ\mathcal{J},\dots$
The Jordan subalgebra $\mathcal{J}$ is called Jordan solvable (resp., Jordan nilpotent) if there exists $n\geq 0$ (resp., $m\geq 0$) such that $\mathcal{J}_{(n)}=0$ (resp., $\mathcal{J}_{m}=0$). The index of a solvable Jordan algebra is defined similarly.
In this paper, we particularly consider the special case where $\mathcal{A}=L_{K}(E)$, $\mathcal{L}=\mathbf{K}_{L_{K}(E)}$ and $\mathcal{J}=\mathbf{S}_{L_{K}(E)}$ with respect to the standard involution on $L_{K}(E)$.
2.3. Homomorphisms of algebras with involutions
Let $K$ be a field, and $(\mathcal{A},^{\star})$ and $(\mathcal{B},^{\natural})$ be $K$-algebras with involutions. A homomorphism of $K$-algebras with involution $\varphi:(\mathcal{A},^{\star})\to(\mathcal{B},^{\natural})$ is a $K$-algebra homomorphism $\varphi:\mathcal{A}\to\mathcal{B}$ such that $(\varphi(ab))^{\natural}=\varphi(b^{\star})\varphi(a^{\star})$.
It is obvious that if $(\mathcal{A},^{\star})$ is isomorphic to $(\mathcal{B},^{\natural})$ as $K$-algebras with involution, then the Lie subalgebras $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{K}_{\mathcal{B}}$, with respect to the involutions and $\natural$ respectively, are also isomorphic as Lie algebras over $K$. Similarly, the Jordan subalgebras $\mathbf{S}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{B}}$ of $\mathcal{A}^{\circ}$ and $\mathcal{B}^{\circ}$, with respect to the involutions and $\natural$ respectively, are also isomorphic as Jordan algebras over $K$.
3. Lie Solvability and Jordan Solvability of $\mathbb{M}_{n}(\mathcal{A})$
Let $K$ be a field, and $\mathcal{A}$ a unital $K$-algebra with involution . Let $\mathbb{M}_{n}(\mathcal{A})$ be the matrix ring of degree $n\geq 1$ over $\mathcal{A}$. Then, it is easy to check that is extended to an involution, which is again denoted by , on $\mathbb{M}_{n}(\mathcal{A})$ by letting
(3.1) $(a_{ij})_{n\times n}^{\natural}=(a_{ji}^{\natural})_{n\times n}$
for every matrix $(a_{ij})\in\mathbb{M}_{n}(\mathcal{A})$. Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be the Lie subalgebra of $\mathbb{M}_{n}(\mathcal{A})^{-}$ with respect to the involution $\natural$ on $\mathbb{M}_{n}(\mathcal{A})$ as defined in (3.1). In this section, we give conditions on $n$ and $\mathcal{A}$ for $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ to be Lie solvable. First, we fix the following notation: In $\mathbb{M}_{n}(\mathcal{A})$, we denote by $E_{ij}$ the matrix whose $(i,j)$-entry is $1$ and the other entries are 0. The following proposition is the main result of this section:
Proposition 3.1.
Let $K$ be a field, and $\mathcal{A}$ be a unital $K$-algebra with an involution . Assume further that $\mathcal{A}$ is an integral domain. Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be as defined above. Then, the following assertions hold:
  1. (a)
    $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ is not Lie solvable when $n\geq 3$.
  2. (b)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie nilpotent.
  3. (c)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable, with index at most $3$. If moreover, the involution on $\mathcal{A}$ is trivial, then the index is $2$.
  4. (d)
    Assume that ${\rm char}(K)\neq 2$. If the involution on $\mathcal{A}$ is trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$; while if the involution on $\mathcal{A}$ is non-trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie solvable.
Proof.
(a): Assume first that $n\geq 3$. Let $a,b,c\in K$ be such that $c\neq 0$ and $a^{2}+b^{2}\neq 0$. We construct inductively a family of matrices $\{A_{m},B_{m},X_{m}\;|\;m=1,2,\dots\}$ in $\mathbb{M}_{n}(\mathcal{A})$ as follows: Put
$A_{1}=a(E_{12}-E_{21})+b(E_{13}-E_{31}),~~B_{1}=c(E_{23}-E_{32}),~~X_{1}=[A_{1},B_{1}];$
and for each $m\geq 2$, we put
$A_{m}=[X_{m-1},B_{m-1}],~~B_{m}=[X_{m-1},A_{m-1}],\text{ and }X_{m}=[A_{m},B_{m}].$
Since $A_{1},B_{1}\in\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$, we get that $X_{1}\in\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}^{(1)}$. Inductively, it can be checked that $A_{m},B_{m}\in\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}^{(m-1)}$, from which it follows that $X_{m}\in\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}^{(m)}$. For each $m\geq 1$, we claim that there exist $a_{m},b_{m},c_{m}\in K$ such that $c_{m}\neq 0$, $a_{m}^{2}+b_{m}^{2}\neq 0$, and
$A_{m}=a_{m}(E_{12}-E_{21})+b_{m}(E_{13}-E_{31}),$
$B_{m}=c_{m}(E_{23}-E_{32}),$
$X_{m}\neq 0.$
The argument proceeds by induction on $m$. For $m=1$, then we set $a_{1}=a$, $b_{1}=b$ and $c_{1}=c$. Then, we have
$A_{1}=a_{1}(E_{12}-E_{21})+b_{1}(E_{13}-E_{31}),$
$B_{1}=c_{1}(E_{23}-E_{32}),$
$X_{1}=-b_{1}c_{1}(E_{12}-E_{21})+a_{1}c_{1}(E_{13}-E_{31})\neq 0.$
Thus, the claim is true for $m=1$. Now, assume that the claim is true for all $m\geq i$ for some $i\geq 2$. We will prove the claim is also true for $m=i+1$. Now, the inductive assumption shows that there exist $a_{i},b_{i},c_{i}\in K$ such that $c_{i}\neq 0,a_{i}^{2}+b_{i}^{2}\neq 0$, and
$A_{i}=a_{i}(E_{12}-E_{21})+b_{i}(E_{13}-E_{31}),$
$B_{i}=c_{i}(E_{23}-E_{32}),$
$X_{i}=[A_{i},B_{i}]=-b_{i}c_{i}(E_{12}-E_{21})+a_{i}c_{i}(E_{13}-E_{31})\neq 0.$
From this, a direct calculation shows that
$A_{i+1}=[X_{i},B_{i}]=-a_{i}c_{i}^{2}(E_{12}-E_{21})-b_{i}c_{i}^{2}(E_{13}-E_{31}),$
$B_{i+1}=[X_{i},A_{i}]=(a_{i}^{2}+b_{i}^{2})c_{i}(E_{23}-E_{32}),$
$X_{i+1}=[A_{i+1},B_{i+1}]=b_{i}(a_{i}^{2}+b_{i}^{2})c_{i}^{3}(E_{12}-E_{21})-a_{i}(a_{i}^{2}+b_{i}^{2})c_{i}^{3}(E_{13}-E_{31}).$
Since $a_{i}^{2}+b_{i}^{2}\neq 0$, we have $X_{i+1}\neq 0$. The claim is proved. It follows that $X_{m}\neq 0$ for all $m\geq 1$, and so $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}^{(m)}\neq 0$ for all $m\geq 1$. Therefore $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is not Lie solvable, proving (a).
To prove the remaining assertions, we assume from now on that $n=2$. Let us calculate $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ more explicitly. It is obvious that
$\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}=\left\{\begin{pmatrix}a&b\\ -b^{\natural}&c\end{pmatrix}\;|\;a,b,c\in\mathcal{A}\text{ and }a=-a^{\natural},c=-c^{\natural}\right\}.$
Let $A_{i},B_{i}$ ($1\leq i\leq 4$) be arbitrary matrices in $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$. Then, these matrices can be written as
$A_{i}=\left(\begin{array}[]{cc}a_{i}&b_{i}\\ -b_{i}^{\natural}&c_{i}\end{array}\right)\text{ and }B_{i}=\left(\begin{array}[]{cc}u_{i}&v_{i}\\ -v_{i}^{\natural}&w_{i}\end{array}\right),$
where $a_{i},b_{i},c_{i},u_{i},v_{i},w_{i}\in\mathcal{A}$ and $a_{i}=-a_{i}^{\natural}$, $c_{i}=-c_{i}^{\natural}$, $u_{i}=-u_{i}^{\natural}$, $w_{i}=-w_{i}^{\natural}$.
For each $1\leq i\leq 4$, put $X_{i}=[A_{i},B_{i}]$. Because $a_{i}=-a_{i}^{\natural}$, $c_{i}=-c_{i}^{\natural}$, $u_{i}=-u_{i}^{\natural}$, and $w_{i}=-w_{i}^{\natural}$, we get
(3.2) $X_{i}=\left(\begin{array}[]{cc}r_{i}&s_{i}\\ -s_{i}^{\natural}&-r_{i}\end{array}\right),\text{ where }$
(3.3) $r_{i}=b_{i}^{\natural}v_{i}-b_{i}v_{i}^{\natural}\text{ and }s_{i}=v_{i}(a_{i}-c_{i})+b_{i}(w_{i}-u_{i}).$
It follows that
(3.4) $X_{1},X_{2}]=\left(\begin{array}[]{cc}s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}&0\\[8.5359pt] 2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)&s_{2}^{\natural}s_{1}-s_{1}^{\natural}s_{2}\end{array}\right),\text{ where }$
(3.5) $s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}\hskip 2.84544pt=(v_{2}v_{1}^{\natural}-v_{1}v_{2}^{\natural})(a_{1}-c_{1})(c_{2}-a_{2})$
$\hskip 54.06006pt+(v_{2}b_{1}^{\natural}-b_{1}v_{2}^{\natural})(a_{2}-c_{2})(u_{1}-w_{1})$
$\hskip 54.06006pt+(v_{1}b_{2}^{\natural}-b_{2}v_{1}^{\natural})(w_{2}-u_{2})(a_{1}-c_{1})$
$\hskip 54.06006pt+(b_{1}b_{2}^{\natural}-b_{2}b_{1}^{\natural})(u_{1}-w_{1})(u_{2}-w_{2}).$
(b) and (c): Assume that ${\rm char}(K)=2$. As $2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)=0$, it follows from (3.4) that $X_{1},X_{2}]$ is a diagonal matrix. Similarly, we also obtain that $X_{3},X_{4}]$ is also a diagonal matrix. Thus $X_{1},X_{2}],[X_{3},X_{4}]]=0$, which means that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(3)}=0$. We divide our situation into two possible cases:
Case 1. ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is trivial. As $b_{i}^{\natural}=b_{i}$ and $v_{i}^{\natural}=v_{i}$, we get from (3.3) that $r_{i}=0$. It follows from (3.2) that $X_{i}=\left(\begin{array}[]{cc}0&s_{i}\\ -s_{i}^{\natural}&0\end{array}\right)$, and so $X_{1},X_{2}]=0$; or equivalently, we have $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(2)}=0$. In other words, $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 2. This completes the proof of the second assertion of (c).
Case 2. ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is non-trivial. In this case, suitable choices of the matrices $A_{1},A_{2},B_{1},B_{2}$ yield that $X_{1},X_{2}]\neq 0$. For example, if we choose $a_{1}=c_{1}$, $a_{2}=c_{2}$, $b_{1}=1$, $u_{1}\neq w_{1}$, $w_{1}=w_{2}=0$, $0\neq u_{1}=u_{2}\in K$, $b_{2}\neq b_{2}^{\natural}$, then it follows from (3.5) that $s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}=(b_{2}^{\natural}-b_{2})u_{1}^{2}\neq 0$. Therefore, we get from (3.4) that $X_{1},X_{2}]\neq 0$. It follows that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 3. This completes the proof of (c).
Now, we give the proof of (b). Let $A=E_{12}+E_{21}$ and $B=E_{11}$, which belong to $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$. Define
[…]
(d): Recall that
$\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}=\left\{\begin{pmatrix}a&b\\ -b^{\natural}&c\end{pmatrix}\;|\;a,b,c\in\mathcal{A}\text{ and }a=-a^{\natural},c=-c^{\natural}\right\}.$
We consider the following two possible cases:
Case 3. ${\rm char}(K)\neq 2$ and the involution on $\mathcal{A}$ is trivial. Because $a^{\natural}=a$, the equation $a=-a^{\natural}$ is equivalent to $2a=0$, and hence $a=0$. Similarly, we also obtain that $c=0$. Moreover, as $b^{\natural}=b$, we finally get that
$\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}=\left\{\begin{pmatrix}0&b\\ -b&0\end{pmatrix}\;|\;b\in\mathcal{A}\right\}.$
It follows that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$.
Case 4. ${\rm char}(K)\neq 2$ and the involution on $\mathcal{A}$ is non-trivial. Fix $a\in\mathcal{A}$ such that $a\neq a^{\natural}$. Put $u=a-a^{\natural}$. Then, we have $u=-u^{\natural}$. We construct inductively a family of matrices $A_{m}$, $B_{m}$, and $X_{m}$ as follows. For $m=1$, we put
$A_{1}=\left(\begin{array}[]{cc}0&u\\ u&0\end{array}\right),~B_{1}=\left(\begin{array}[]{cc}u&0\\ 0&-u\end{array}\right)\text{ and }X_{1}=[A_{1},B_{1}],$
For $m\geq 2$, we put
$A_{m}=[X_{m-1},B_{m-1}],B_{m}=[X_{m-1},A_{m-1}]\text{ and }X_{m}=[A_{m},B_{m}].$
Since $A_{1},B_{1}\in\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$, we get that $X_{1}\in\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(1)}$. Inductively, it can be checked that $A_{m},B_{m}\in\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(m-1)}$, from which it follows that $X_{m}\in\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(m)}$. We claim that for each positive integer $m$, there exists $v_{m}\in\mathcal{A}$ such that $0\neq v_{m}=-v_{m}^{\natural}$ and
$A_{m}=\left(\begin{array}[]{cc}0&v_{m}\\ v_{m}&0\end{array}\right),$
$B_{m}=\left(\begin{array}[]{cc}(-1)^{m+1}v_{m}&0\\ 0&(-1)^{m}v_{m}\end{array}\right),$
$X_{m}=\left(\begin{array}[]{cc}0&(-1)^{m}2v_{m}^{2}\\ (-1)^{m+1}2v_{m}^{2}&0\end{array}\right)\neq 0.$
We proceed by induction on $m$. For $m=1$, we set $v_{1}=u$. Then, we have $0\neq v_{1}=v_{1}^{\natural}$ and $X_{1}=[A_{1},B_{1}]=\left(\begin{array}[]{cc}0&-2u^{2}\\ 2u^{2}&0\end{array}\right)\neq 0$. Thus, the claim is true for the case $m=1$. Now, assume that the claim is true for $m=i\geq 2$. That means there exist $v_{i}\in\mathcal{A}$ such that $0\neq v_{i}=-v_{i}^{\natural}$ and
$A_{i}=\left(\begin{array}[]{cc}0&v_{i}\\ v_{i}&0\end{array}\right),$
$B_{i}=\left(\begin{array}[]{cc}(-1)^{i+1}v_{i}&0\\ 0&(-1)^{i}v_{i}\end{array}\right),$
$X_{i}=\left(\begin{array}[]{cc}0&(-1)^{i}2v_{i}^{2}\\ (-1)^{i+1}2v_{i}^{2}&0\end{array}\right)\neq 0.$
Now, we prove the claim for the case $m=i+1$. A straightforward calculation shows that
$A_{i+1}=\left(\begin{array}[]{cc}0&4v_{i}^{3}\\ 4v_{i}^{3}&0\end{array}\right),$
$B_{i+1}=\left(\begin{array}[]{cc}(-1)^{i}4v_{i}^{3}&0\\ 0&(-1)^{i+1}4v_{i}^{3}\end{array}\right),$
$X_{i+1}=\left(\begin{array}[]{cc}0&(-1)^{i+1}32v_{i}^{6}\\ (-1)^{i+2}32v_{i}^{6}&0\end{array}\right)\neq 0.$
Now, put $v_{i+1}=4v_{i}^{3}$. Then, we have
$v_{i+1}^{\natural}=(4v_{i}^{3})^{\natural}=4(v_{i}^{\natural})^{3}=-4v_{i}^{3}=-v_{i+1}\neq 0.$
Therefore, the claim is true for $m=i+1$. The claim is proved. It follows that $X_{m}\neq 0$ for all $m\geq 1$. In other words, we have $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(m)}$ for all $m\geq 1$, which means that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie solvable. ∎
In the remainder of this section, we record two corollaries of Proposition 3.1 for further use. In the proposition, let $A=K$ and be the trivial involution on $K$. Then, the extended involution on $\mathbb{M}_{n}(K)$ given by (3.1) is just the transpose operation on $\mathbb{M}_{n}(K)$; that is, for every matrix $(a_{ij})_{n\times n}\in\mathbb{M}_{n}(K)$, we have
(3.6) $\left((a_{ij})_{n\times n}\right)^{\natural}=\left(a_{ji}\right)_{n\times n}.$
With respect to this involution, the Lie algebra $\mathbf{K}_{\mathbb{M}_{n}(K)}$ consists precisely of skew-symmetric matrices in $\mathbb{M}_{n}(K)$:
$\mathbf{K}_{\mathbb{M}_{n}(K)}=\left\{(a_{ij}):a_{ji}=-a_{ij}\text{ for all }1\leq i,j\leq n\right\}.$
As a consequence of Proposition 3.1, we easily obtain the following:
Corollary 3.2.
Let $K$ be a field, and $n\geq 2$ an integer. Let $\mathbf{K}_{\mathbb{M}_{n}(K)}$ be the Lie subalgebra of skew-symmetric matrices of $\mathbb{M}_{n}(K)^{-}$. The following assertions hold:
  1. (1)
    Suppose that $n=2$. If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)},\mathbf{K}_{\mathbb{M}_{n}(K)}]=0$; and, if ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is Lie solvable of index $2$ but not Lie nilpotent.
  2. (2)
    If $n\geq 3$, then $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is not Lie solvable.
For any field $K$, let $K[x,x^{-1}]$ be the Laurent polynomial $K$-algebra in $x$; that is,
$K[x,x^{-1}]=\left\{\sum_{i=k}^{n}a_{i}x^{i}\;|\;a_{i}\in K\text{ and }k\geq n\in\mathbb{Z}\right\}.$
Then, the map $^{\natural}:K[x,x^{-1}]\to K[x,x^{-1}]$ defined by $(\sum_{i}a_{i}x^{i})^{\natural}=\sum_{i}a_{i}x^{-i}$ is a non-trivial involution on $K[x,x^{-1}]$. Now, in Proposition 3.1, we consider the case where $\mathcal{A}=K[x,x^{-1}]$ and extend the involution on $K[x,x^{-1}]$ to the matrix ring $\mathbb{M}_{n}(K[x,x^{-1}])$ as defined in (3.1). This extended involution $^{\natural}:\mathbb{M}_{n}(K[x,x^{-1}])\to\mathbb{M}_{n}(K[x,x^{-1}])$ is precisely given by the following rule:
(3.7) $\left((f_{ij}(x))_{n\times n}\right)^{\natural}=\left(f_{ji}(x^{-1})\right)_{n\times n}$
for each matrix $(f_{ij}(x))\in\mathbb{M}_{n}(K[x,x^{-1}])$. With respect to this involution, we easily get the following corollary to Proposition 3.1:
Corollary 3.3.
Let $K[x,x^{-1}]$ be the Laurent polynomial $K$-algebra. Let $\mathbf{K}_{\mathbb{M}_{n}(K[x,x^{-1}])}$ be the Lie subalgebra of skew-symmetric elements of $\mathbb{M}_{n}(K[x,x^{-1}])^{-}$ with respect to the involution defined by ((3.7)). Then, the following assertions hold:
  1. (1)
    If $\mathrm{char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{2}(K[x,x^{-1}])}$ is Lie solvable of index $3$ but not Lie nilpotent.
  2. (2)
    If $\mathrm{char}(K)\neq 2$, then $\mathbf{K}_{\mathbb{M}_{2}(K[x,x^{-1}])}$ is not Lie solvable.
4. Lie Solvability and Jordan Solvability of $L_{K}(E)$
Throughout this section, we will denote by $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ the Lie subalgebra of $L_{K}(E)^{-}$ and the Jordan subalgebra of $L_{K}(E)^{\circ}$with respect to standard involution on $L_{K}(E)$ as defined before, respectively. The main aim of this section is to provide a necessary and sufficient condition on the graph $E$ and the field $K$ for $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ to be Lie and Jordan solvable, respectively. We start this section with a lemma which determines a $K$-base of $L_{K}(E)$:
Lemma 4.1 ([1, Corollary 1.5.12]).
Let $E$ be a graph and $K$ a field. Put $\mathscr{A}=\{\lambda\nu^{*}\;|\;\lambda,\nu\in{\rm Path}(E)\text{ and }r(\lambda)=r(\nu)\}$. For each $v\in{\rm Reg}(E)$, let $\{e_{1}^{v},\dots,e_{n_{v}}^{v}\}$ be an enumeration of the elements of $s^{-1}(v)$. Then, a $K$-basis of $L_{K}(E)$ is given by the family
$\mathscr{B}=\mathscr{A}\backslash\{\lambda e_{n_{v}}^{v}(e_{n_{v}}^{v})^{*}\nu^{*}\;:\;r(\lambda)=r(\nu)=v\in{\rm Reg}(E)\}.$
The next lemma provides a necessary condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. The ideas used to prove this lemma come from the proof of [12, Lemma 2.2 and Theorem 2.3]. It can be said that, in $L_{K}(E)$, we can easily choose a suitable subalgebra which is isomorphic to the matrix ring $\mathbb{M}_{3}(K)$. Then, Corollary 3.2 is applied to get all possible cases for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. For convenience of the readers, we still include the proof here.
Lemma 4.2.
Let $L_{K}(E)$ be the Leavitt path algebra of a graph $E$ with coefficients in a field $K$. Then, $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable if one of the following cases occurs:
  1. (1)
    $E$ contains a cycle with an exit.
  2. (2)
    $E$ contains one of following subgraphs:
    $F_{1}:\bullet\bullet\bulletvveeff$
    $F_{2}:\bullet\bullet\bulletvveeff$
    $F_{3}:\bullet\bulletvveeff$
Proof.
(1) Assume that $E$ contains a cycle $c$ with an exit $f$. Then, according to Lemma 4.1, the set $\{c^{m}ff^{*}(c^{*})^{n}:1\leq m,n\leq 3\}$ is linearly independent over $K$ satisfying the rule
$(c^{m}ff^{*}(c^{*})^{i})(c^{j}ff^{*}(c^{*})^{n})=\delta_{ij}c^{m}ff^{*}(c^{*})^{n},$
where $\delta_{ij}$ is the Kronecker delta. It follows that the $K$-subalgebra of $L_{K}(E)$, say $A$, generated by $\{c^{m}ff^{*}(c^{*})^{n}:1\leq m,n\leq 3\}$ is isomorphic to $\mathbb{M}_{3}(K)$. The restriction of the standard involution on $L_{K}(E)$ to $A$ is still an involution on $A$ which satisfies $\mathbf{K}_{A}\subseteq\mathbf{K}_{L_{K}(E)}$. Let be the involution on $\mathbb{M}_{3}(K)$ as defined in (3.6). Then, it can be checked that ($A,^{\star}$) is isomorphic to $(\mathbb{M}_{3}(K),^{\natural})$ as $K$-algebras with involution. Thus, we have that $\mathbf{K}_{A}$ is isomorphic to $\mathbf{K}_{\mathbb{M}_{3}(K)}$ as Lie algebras over $K$. Because $\mathbf{K}_{\mathbb{M}_{3}(K)}$ is not Lie solvable (see Corollary 3.2), it follows that $\mathbf{K}_{A}$, and hence $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable.
(2) For the proof of this assertion, we assume first that $E$ contains $F_{1}$. Then, $v,f,ef$ are all paths ending at $v$. It is straightforward to check the subalgebra $B$ of $L_{K}(E)$ generated by $v$, $f$, efef, $f^{*}$, $^{*}e^{*}, ff$$f^{*}$, $^{*}, ffeff^{*}e^{*}$, $eff^{*}e^{*}$ is isomorphic to $\mathbb{M}_{3}(K)$ via map: $eff^{*}e^{*}\mapsto E_{11}$, $ff^{*}\mapsto E_{22}$, $v\mapsto E_{33}$, $eff^{*}\mapsto E_{12}$, $ff^{*}e^{*}\mapsto E_{21}$, $f\mapsto E_{23}$, $f^{*}\mapsto E_{32}$, $ef\mapsto E_{13}$, and $f^{*}e^{*}\mapsto E_{31}$ (see also a similar techniques used in the proof of [1, Lemma 2.6.4]). The restriction of the involution to $B$ is also an involution on $B$ for which $\mathbf{K}_{B}\subseteq\mathbf{K}_{L_{K}(E)}$. Thus, $\mathbf{K}_{B}$ is isomorphic to $\mathbf{K}_{\mathbb{M}_{3}(K)}$ as Lie algebras over $K$. Therefore, by Corollary 3.2, we conclude that $\mathbf{K}_{B}$, and so $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable.
Finally, assume that $E$ contains $F_{2}$ or $F_{3}$. Then, the subalgebra $C$ generated by $v,e,f,e^{*},f^{*},ee^{*},ff^{*},ef^{*},fe^{*}$ is isomorphic to $\mathbb{M}_{3}(K)$ via map $ee^{*}\mapsto E_{11}$, $ff^{*}\mapsto E_{22}$, $v\mapsto E_{33}$, $ef^{*}\mapsto E_{12}$, $fe^{*}\mapsto E_{21}$, $e\mapsto E_{13}$, $e^{*}\mapsto E_{31}$, $f\mapsto E_{23}$, $f^{*}\mapsto E_{32}$. A similar argument as in previous paragraph shows that $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable too. The proof is now completed. ∎
As a consequence of Lemma 4.2, we have the following remark.
Remark 1.
Let $L_{K}(E)$ be the Leavitt path algebra of a graph $E$ with coefficients in a field $K$. If $\mathbf{K}_{L_{K}(E)}$ is Lie solvable then $E$ is a disjoint union of following graphs:
$E_{1}:\bullet$
$E_{2}:\bullet$
$E_{3}:\bullet\bullet$
[…]
in which $\{u_{i}|i\in I\}$ and $\{v_{j}|j\in J\}$, where $I$ and $J$ are indexing non-empty sets of arbitrary cardinality.
Remark 2.
Let $E_{4},E_{5},E_{6}$ be graphs as defined above. If $m:=|I|$ and $n:=|J|$ are finite, then by [1, Corollary 2.7.5], we get that
$L_{K}(E_{1})\cong K,\;\;L_{K}(E_{2})\cong K[x,x^{-1}],\;\;L_{K}(E_{3})\cong\mathbb{M}_{2}(K[x,x^{-1}]),$
$L_{K}(E_{4})\cong\bigoplus_{i=1}^{n}\mathbb{M}_{2}(K),\;\;L_{K}(E_{5})\cong\bigoplus_{i=1}^{m}\mathbb{M}_{2}(K[x,x^{-1}]),$
$L_{K}(E_{6})\cong\bigoplus_{i=1}^{n}\mathbb{M}_{2}(K)\oplus\bigoplus_{i=1}^{m}\mathbb{M}_{2}(K[x,x^{-1}]).$
In what follows, we give a sufficient condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. We proceed by proving several auxiliary lemmas.
Lemma 4.3.
Let $\mathcal{A}$ be a $K$-algebra with an involution $\star$, and $\mathbf{J}$ be an indexing set of arbitrary cardinality. Let $I_{j},j\in\mathbf{J}$, be ideals of $\mathcal{A}$ satisfying the property that for each $j$, if $a\in I_{j}$ then $a^{\star}\in I_{j}$. Put $M=\sum_{j\in\mathbf{J}}I_{j}$. If $M=\bigoplus_{j\in\mathbf{J}}I_{j}$, then $\mathbf{K}_{M}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}$. Consequently, $\mathbf{K}_{M}$ is Lie solvable ((nilpotent)) if and only if $\mathbf{K}_{I_{j}}$ is Lie solvable ((nilpotent)) of a common index for all $j\in\mathbf{J}$.
Proof.
Let $x\in\mathbf{K}_{M}$. Then, we may write $x=a_{j_{1}}+a_{j_{2}}+\dots+a_{j_{n}}$, where $a_{j_{k}}\in I_{j_{k}}$ for all $1\leq k\leq n$. The equation $x^{\star}=-x$ implies that
$a_{j_{1}}+a_{j_{1}}^{\star}=-(a_{j_{2}}+a_{j_{2}}^{\star}+\cdots+a_{j_{n}}+a_{j_{n}}^{\star}).$
Because $a_{j_{k}}\in I_{j_{k}}$ for all $i$, the above equation shows that $a_{j_{1}}+a_{j_{1}}^{\star}\in I_{j_{1}}\cap(I_{j_{2}}\oplus\cdots\oplus I_{j_{n}})=0$. It follows that $a_{1}+a_{1}^{\star}=0$, and so $a_{j_{1}}\in\mathbf{K}_{I_{j_{1}}}$. Similarly, we also obtain that $a_{i}\in\mathbf{K}_{I_{j_{k}}}$ for $k\in\{2,\dots,n\}$. In other words, we have $\mathbf{K}_{M}\subseteq\sum_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}$. The inverse inclusion is clear. The final assertion follows from the fact that $\mathbf{K}_{M}^{(n)}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}^{(n)}$ and $\mathbf{K}_{M}^{n}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}^{n}$ for all integer $n\geq 1$. ∎
Lemma 4.4.
Let $K$ be a field, and $E_{1},E_{2},E_{3}$ be graphs as defined in Remark 1. Then, the following assertions hold:
  1. (1)
    $\mathbf{K}_{L_{K}(E_{1})}$ and $\mathbf{K}_{L_{K}(E_{2})}$ are Lie solvable of index $1$.
  2. (2)
    $\mathbf{K}_{L_{K}(E_{3})}$ is Lie solvable ((of index 33)) if and only if $\mathrm{char}(K)=2$. Additionally, $\mathbf{K}_{L_{K}(E_{3})}$ is not Lie nilpotent.
Proof.
The assertion (i) follows immediately from the fact that $L_{K}(E_{1})$ and $L_{K}(E_{2})$ are commutative. Moreover, because $L_{K}(E_{3})\cong\mathbb{M}_{2}(K[x,x^{-1}])$, we get that $\mathbf{K}_{L_{K}(E_{3})}$ is not Lie nilpotent. ∎
Lemma 4.5.
Let $K$ be a field, and $E_{4},E_{5},E_{6}$ be graphs as defined in Remark 1. Then, the following assertions hold:
  1. (1)
    If $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable, then it is Lie solvable of index $l\leq 3$. More precisely, if $|I|$ is finite and ${\rm char}(K)\neq 2$ ((resp., ${\rm char}(K)=2$)), then $l=1$ ((resp., $l=2$)). If $|I|$ is infinite and ${\rm char}(K)\neq 2$ ((reps., ${\rm char}(K)=2$)), then $l=2$ ((resp., $l=3$)). Additionally, $\mathbf{K}_{L_{K}(E_{4})}$ is Lie nilpotent if and only if $|I|$ is finite and ${\rm char}(K)\neq 2$, if and only if $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]=1$.
  2. (2)
    $\mathbf{K}_{L_{K}(E_{5})}$ is Lie solvable ((of index $l\leq 4$)) if and only if ${\rm char}(K)=2$. More precisely, if $|J|$ is finite ((resp., infinite)), then $l=3$ ((resp., $l=4$)). Additionally, $\mathbf{K}_{L_{K}(E_{5})}$ is not Lie nilpotent.
  3. (3)
    $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable ((of index $l\leq 4$)) if and only if ${\rm char}(K)=2$. More precisely, if both $|I|$ and $|J|$ are finite ((resp., either $|I|$ or $|J|$ is infinite)), then $l=3$ ((resp., $l=4$)). Additionally, $\mathbf{K}_{L_{K}(E_{6})}$ is not Lie nilpotent.
Proof.
(1) Consider the graph $E_{4}$. For each $1\leq i\leq n$, let $I_{i}$ be the ideal of $L_{K}(E_{4})$ generated by $u_{i}$. Then, it is clear that, for each $i$,
$I_{i}={\rm span}_{K}\{u_{i},e_{i},e_{i}^{*},e_{i}e_{i}^{*}\}.$
Moreover, according to Lemma 4.1, the set $\{u_{i},e_{i},e_{i}^{*},e_{i}e_{i}^{*}\}$ is linearly independent over $K$. Thus, it is can be checked that $I_{i}\cong\mathbb{M}_{2}(K)$ via the mapping $u_{i}\mapsto E_{11}$, $e_{i}e_{i}^{*}\mapsto E_{22}$, $e_{i}\mapsto E_{12}$ and $e_{i}^{*}\mapsto E_{21}$. Moreover, we have $\sum_{i=1}^{n}I_{i}=\bigoplus_{i=1}^{n}I_{i}$. Put $M=\bigoplus_{i=1}^{n}I_{i}$. Then, according to Lemma 4.3 and Corollary 3.2, we conclude that $\mathbf{K}_{M}=\bigoplus_{i=1}^{n}\mathbf{K}_{I_{i}}$, and that $\mathbf{K}_{M}$ is Lie solvable. Moreover, in view of Corollary 3.2, if ${\rm char}(K)\neq 2$, then $\mathbf{K}_{M},\mathbf{K}_{M}]=0$; and, if ${\rm char}(K)=2$ then $\mathbf{K}_{M}$ is Lie solvable of index $2$ but not Lie nilpotent. Now, we consider two possible cases:
Case 1. $n$ is finite. In this case, as $u\in M$, we get that $L_{K}(E_{4})=M$, from which it follows that $\mathbf{K}_{L_{K}(E_{4})}=\mathbf{K}_{M}$. Thus, all conclusions follow immediately in this case.
Case 2. $n$ is infinite. In this case, $u$ is an infinite emitter and $u\not\in M$. This implies that $L_{K}(E_{4})=Ku\oplus M$ as $K$-vector spaces. It follows that
$\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq[L_{K}(E_{4}),L_{K}(E_{4})]=[M,M]\subseteq M.$
As $M$ is Lie solvable, we conclude that $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$, and hence $\mathbf{K}_{L_{K}(E_{4})}$, is Lie solvable. To determine the solvable index of $\mathbf{K}_{L_{K}(E_{4})}$, we consider the following subcases:
[…]
Subcase 2.2. ${\rm char}(K)=2$. As $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq\mathbf{K}_{M}$ which is Lie solvable of index $2$, it follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index at most $3$. Because $v\in\mathbf{K}_{L_{K}(E_{4})}\backslash\mathbf{K}_{M}$, we conclude that $\mathbf{K}_{L_{K}(E_{4})}\neq[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$. It follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index $3$. On the other hand, as $\mathbf{K}_{I_{i}}$ is not Lie nilpotent and $\mathbf{K}_{I_{i}}\subseteq\mathbf{K}_{L_{K}(E_{4})}$ for all $i$, we obtain that $\mathbf{K}_{L_{K}(E_{4})}$ is not Lie nilpotent. The proof of assertion (1) is complete.
(2) Consider the graph $E_{5}$. For each $1\leq i\leq m$, let $J_{i}$ be the ideal of $L_{K}(E_{5})$ generated by the $v_{j}$’s. Then, we have $J_{j}\cong\mathbb{M}_{2}(K[x,x^{-1}])$ for all $j$, and $\sum_{i=1}^{n}J_{j}=\bigoplus_{j=1}^{n}J_{j}$. Therefore, one can repeat the arguments of the proof of assertion(1), together with using Corollary 3.3 instead of Corollary 3.2, to get (2).
(3) Consider the graph $E_{6}$. Let $I_{i}$ and $J_{j}$ be respectively the ideals of $L_{K}(E_{6})$ generated by $v_{i}$ and $u_{j}$ for $1\leq i\leq n$ and $1\leq j\leq m$. Then, for each $i$ and each $j$, we have $I_{i}\cong\mathbb{M}_{2}(K)$ and $J_{j}\cong\mathbb{M}_{2}(K[x,x^{-1}])$. It can be checked that
$\sum_{i=1}^{n}I_{i}+\sum_{j=1}^{m}J_{j}=\bigoplus_{i=1}^{n}I_{i}\oplus\bigoplus_{j=1}^{m}J_{j}.$
Put $M=\bigoplus_{i=1}^{n}I_{i}$ and $N=\bigoplus_{j=1}^{m}J_{j}$. Then, as before, $\mathbf{K}_{M}$ is Lie solvable of index at most $3$; and, $\mathbf{K}_{N}$ is Lie solvable (of index at most $4$) if and only if ${\rm char}(K)=2$. Moreover, it is clear that $L_{K}(E_{6})=Kw+(M\oplus N)$. As $N$ is not Lie nilpotent, we get that $\mathbf{K}_{L_{K}(E_{6})}$ is not Lie nilpotent. Moreover, we have that $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable if and only if both $\mathbf{K}_{M}$ and $\mathbf{K}_{N}$ are Lie solvable, if and only if ${\rm char}(K)=2$. Now, assume that ${\rm char}(K)=2$. To determine the Lie solvability index of $\mathbf{K}_{L_{K}(E_{6})}$, we consider the following cases. If both $|I|$ and $|J|$ are finite, then $w\in M\oplus N$; and so, $L_{K}(E_{6})=M\oplus N$. Again, by Lemma 4.3, we obtain that $\mathbf{K}_{L_{K}(E_{6})}=\mathbf{K}_{M}\oplus\mathbf{K}_{N}$. According to Corollary 3.2, $\mathbf{K}_{M}$ is Lie solvable of index $2$, while by Corollary 3.3, $\mathbf{K}_{N}$ is Lie solvable of index $3$. It follows that $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable of index $3$. If either $|I|$ or $|J|$ is infinite, then $w\notin M\oplus N$. Thus, a similar argument as used in Case 2 above shows that $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable of index $4$. ∎
We now present the main result of this section: a sufficient and necessary condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable.
Theorem 4.6.
Let $K$ be a field and $E$ a graph. Then, the followings hold:
  1. (1)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 4$)) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$. Specially, the index is $4$ precisely when $E$ contains either $E_{5}$ or $E_{6}$ with infinite $|I|$ or $|J|$.
  2. (2)
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 2$)) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$. Specially, the index is $2$ precisely when $E$ contains $E_{4}$ with infinite $|I|$.
Proof.
To prove (1), we assume first that ${\rm char}(K)=2$. For a proof of the “only if” part, we assume that $L_{K}(E)$ is Lie solvable. Then by Remark 1, the graph $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$. Conversely, we assume that $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$. Then, $L_{K}(E)$ is a direct summand of ideals each of which is isomorphic to $L_{K}(E_{1})$, $\dots$, $L_{K}(E_{6})$. In view of Lemmas 4.4 and 4.5, we get that $L_{K}(E_{1})$, $\dots$, $L_{K}(E_{6})$ are Lie solvable, and so the assertion (1) follows. Finally, the proof of (2) is similar to that of (1). ∎
If $E$ is row-finite, then by Remark 2, we can obtain more detailed descriptions as follows:
Corollary 4.7.
Let $K$ be a field and $E$ a row-finite graph.
  1. (1)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 3$)) if and only if $E$ is a disjoint union graphs $E_{1},\dots,E_{6}$ with finite $|I|$ and $|J|$. In this case,
    $L_{K}(E)\cong\bigoplus_{I_{1}}K\oplus\bigoplus_{I_{2}}K[x,x^{-1}]\oplus\bigoplus_{I_{3}}\mathbb{M}_{2}(K)\oplus\bigoplus_{I_{4}}\mathbb{M}_{2}(K[x,x^{-1}]),$
    for suitable ((possibly infinite)) indexing sets $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$. Moreover, the index is $3$ precisely when $E$ contains $E_{3}$, $E_{5}$ or $E_{6}$.
  2. (2)
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 2$)) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$ with finite $|I|$. In this case, we have
    $L_{K}(E)\cong\bigoplus_{J_{1}}K\oplus\bigoplus_{J_{2}}K[x,x^{-1}]\oplus\bigoplus_{J_{3}}\mathbb{M}_{2}(K),$
    for suitable ((possibly infinite)) indexing sets $J_{1}$, $J_{2}$, $J_{3}$. Moreover, the index is $2$ precisely when $E$ contains $E_{3}$.
The following theorem provides a necessary and sufficient condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie nilpotent. The proof of this theorem is similar to that of Theorem 4.6, so it should be omitted.
Theorem 4.8.
Let $K$ be a field and $E$ a graph.
  1. (1)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $E$ is a disjoint union graphs $E_{1}$ and $E_{2}$.
  2. (2)
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $E$ is a disjoint union graphs $E_{1}$, $E_{2}$, and $E_{4}$ with finite $n$.
In particular, $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $\mathbf{K}_{L_{K}(E)},\mathbf{K}_{L_{K}(E)}]=0$.
Corollary 4.9.
Let $K$ be a field and $E$ a graph. If $E$ contains an infinite emitter, then $\mathbf{K}_{L_{K}(E)}$ is not Lie nilpotent.
The following corollary addresses a situation where being Lie solvable and being Lie nilpotent are equivalent.
Corollary 4.10.
Let $K$ be a field and $E$ a row-finite graph. If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable if and only if it is Lie nilpotent, if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$.
Proof.
Assume that $L_{K}(E)$ is Lie solvable. In view of Corollary 4.7, $E$ is a disjoint union of a finite number of graphs $E_{1}$, $E_{2}$ or $E_{4}$ with finite $|I|$. Therefore, $L_{K}(E)$ is Lie nilpotent by Theorem 4.8. ∎
In the remainder of this section, we investigate the solvability of the Jordan algebra $\mathbf{S}_{L_{K}(E)}$.
Theorem 4.11.
Let $K$ be a field and $E$ a graph. Then, the following assertions hold:
  1. (1)
    If ${\rm char}(K)=2$, then $\mathbf{S}_{L_{K}(E)}$ is Jordan solvable ((resp., nilpotent)) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$ ((resp., $E_{1}$ and $E_{2}$)).
  2. (2)
    If ${\rm char}(K)\neq 2$, then $\mathbf{S}_{L_{K}(E)}$ is not Jordan nilpotent.
Proof.
(1) Assume that $\mathrm{char}(K)=2$. Then,
$a\circ b=ab+ba=ab-ba=[a,b]\text{ for any }a,b\in L_{K}(E).$
It follows that $\mathbf{S}_{L_{K}(E)}=\mathbf{K}_{L_{K}(E)}$. Thus, $\mathbf{S}_{L_{K}(E)}$ is Jordan solvable (resp., nilpotent) if and only if $\mathbf{K}_{L_{K}(E)}$ is Lie solvable (resp., nilpotent). By Theorems 4.6 and 4.8, all conclusions of (1) are asserted.
(2) Assume that $\mathrm{char}(K)\neq 2$. Let $v$ be an arbitrary vertex of $E$. Then, it is clear that $v\in\mathbf{S}_{L_{K}(E)}$. Put $X_{0}=v$ and $X_{m+1}=X_{m}\circ v$, for all $m\geq 0$. Then, it can be checked that $X_{m}=2^{m}v$ for all $m\geq 0$. Since $\mathrm{char}(K)\neq 2$, it follows that $X_{m}\neq 0$ for all $m\geq 0$. Hence, $\mathbf{S}_{L_{K}(E)}$ is not Jordan nilpotent. ∎
To conclude this paper, we relate the solvability of the four structures $L_{K}(E)^{\circ}$, $\mathbf{S}_{L_{K}(E)}$, $L_{K}(E)^{-}$ and $\mathbf{K}_{L_{K}(E)}$.
Remark 3.
Theorem 4.11 is also true if we replace $\mathbf{S}_{L_{K}(E)}$ by the whole $L_{K}(E)^{\circ}$. This means $L_{K}(E)^{\circ}$ is Jordan solvable or Jordan nilpotent if and only if $\mathbf{S}_{L_{K}(E)}$ is. At the other extreme, in view of Theorem 4.6(2), if ${\rm char}(K)\neq 2$ then $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable but $L_{K}(E_{4})^{-}$ is not Lie solvable. These facts show that $\mathbf{K}_{L_{K}(E)}$ is Lie solvable for a wider class of graph $E$ than $L_{K}(E)^{-}$. Also, in some senses, the Lie algebra $\mathbf{S}_{L_{K}(E)}$ is bigger than the Jordan algebra $\mathbf{K}_{L_{K}(E)}$ in $L_{K}(E)$.
Funding. This work is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under Grant No. 101.04-2025.41.